# EquiReg: Equivariance Regularized Diffusion for Inverse Problems

### v1

@@ Abstract
Diffusion models represent the state-of-the-art for solving inverse problems such as image restoration tasks. In the Bayesian framework, diffusion-based inverse solvers incorporate a likelihood term to guide the prior sampling process, generating data consistent with the posterior distribution. However, due to the intractability of the likelihood term, many current methods rely on isotropic Gaussian approximations, which lead to deviations from the data manifold and result in inconsistent, unstable reconstructions. We propose Equivariance Regularized (EquiReg) diffusion, a general framework for regularizing posterior sampling in diffusion-based inverse problem solvers. EquiReg enhances reconstructions by reweighting diffusion trajectories and penalizing those that deviate from the data manifold. We define a new distribution-dependent equivariance error, empirically identify functions that exhibit low error for on-manifold samples and higher error for off-manifold samples, and leverage these functions to regularize the diffusion sampling process. When applied to a variety of solvers, EquiReg outperforms state-of-the-art diffusion models in both linear and nonlinear image restoration tasks, as well as in reconstructing partial differential equations.
Figure 1: Ground truth, measurement, baseline, and reconstruction. R1 (left-to-right): super-resolution and motion deblur. R2: box and random inpainting. R3: PDE solving and phase retrieval.
@@ 1 Introduction
Inverse problems consider the problem of recovering an unknown signal ${\bm{x}}^{*}\in\mathbb{R}^{d}$ from undersampled noisy measurements:
@@
where ${\mathcal{A}}$ is a known measurement operator, and $\bm{\nu}$ is an unknown noise [1]. Inverse problems are widely applied and studied in science and engineering, including imaging science [2], medical imaging [3, 4], image restorations [5, 6, 7, 8], computational neuroscience [9], seismology [10], radar imaging [11, 12], source separation [13, 14, 15], and astrophotography [16].
Figure 2: Equivariance Regularized (EquiReg) diffusion for solving inverse problems. a) EquiReg regularizes the posterior sampling trajectory, addressing the errors introduced due to the posterior mean expectation. b) The EquiReg loss function is lower for on- or near-manifold (clean) data and higher for off- or far-from-manifold (blurred and missing pixel) data.
Figure 3: Illustration of off-manifold posterior expectation. a) Manifold interpretation. While samples from $p({\bm{x}}_{0}|{\bm{x}}_{t})$ may lie on the manifold, their posterior expectation, being a linear combination, can fall off the manifold. b) Distribution interpretation. In a bimodal distribution, the posterior expectation lies in a low probability region.
Inverse problems are ill-posed, i.e., the inversion process can have many solutions; hence, they require prior information about the desired solution [17]. Classical approaches use regularization techniques to improve the stability and offer uniqueness of the inversion, formulated as $\min_{{\bm{x}}\in\mathbb{R}^{d}}\ \|{\bm{y}}-{\mathcal{A}}({\bm{x}})\|^{2}+\mathcal{R}({\bm{x}})$, where the first term encapsulates the forward model error and $\mathcal{R}$ is a function regularizing the solution. Bayesian inference formulates the inverse problem as finding ${\bm{x}}$ maximizing the posterior distribution $p({\bm{x}}|{\bm{y}})\propto p({\bm{y}}|{\bm{x}})p({\bm{x}})$ where $p({\bm{y}}|{\bm{x}})$ is the likelihood of the measurements given the signal, and $p({\bm{x}})$ is a prior characterizing the signal structure [18]. Sparsity [19, 20, 21, 22, 23], low-rank [24, 25], and total variation [26] are examples of hand-crafted regularizers, imposing a prior model.
@@
This paper focuses on methods that leverage unconditionally pre-trained score-based generative diffusion models as learned priors [27, 28, 29, 30]. These methods define a sequential noising process ${\bm{x}}_{0}\sim p_{\mathrm{data}}\to{\bm{x}}_{t}\to{\bm{x}}_{T}\sim p_{T}({\bm{x}})\approx\mathcal{N}(\bm{0},{\bm{I}})$ and a reverse denoising process parameterized by a neural network score $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{x}}_{t})$ [31, 30]. During sampling, these approaches incorporate gradient signals carrying likelihood-related information to solve inverse problems. Applications include image restoration [32, 33, 34, 35, 36, 37, 38, 39, 40, 41], medical imaging [42, 43, 44, 45, 46, 47, 48, 49], and solving partial differential equations (PDEs) [50, 51, 52, 53, 54, 55, 56, 57].
@@
Solving inverse problems with diffusion [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 51, 52, 58, 59] requires computing the conditional score $\nabla_{{\bm{x}}_{t}}\log{p_{t}({\bm{x}}_{t}|{\bm{y}})}$, decomposed into $\nabla_{{\bm{x}}_{t}}\log{p_{t}({\bm{x}}_{t})}+\nabla_{{\bm{x}}_{t}}\log{p_{t}({\bm{y}}|{\bm{x}}_{t})}$. This introduces challenges, as the likelihood score $\nabla_{{\bm{x}}_{t}}\log{p_{t}({\bm{y}}|{\bm{x}}_{t})}$ is only computationally tractable when $t=0$. To handle the intractable likelihood for $t>0$, many state-of-the-art methods approximate the posterior $p_{t}({\bm{x}}_{0}|{\bm{x}}_{t})$ with the isotropic Gaussian distribution [32, 38, 39, 40, 60], where the distribution expectation ${\bm{x}}_{0|t}\coloneqq\mathbb{E}[{\bm{x}}_{0}|{\bm{x}}_{t}]$ is computed via Tweedie’s formula using the optimal denoising score $s_{\theta}({\bm{x}}_{t},t)$.
This Gaussian approximation may be inaccurate (e.g., for a bimodal or a highly complex distribution). Moreover, this poor estimation can result in very high errors in likelihood score computation, especially when point estimations are employed as in DPS and PSLD [32, 38]. Conditional expectation, as in the posterior mean expectation, computes a linear combination of all possible ${\bm{x}}_{0}$; hence, from the manifold perspective, the posterior mean expectation may end up being located off-manifold (see Figure 3). The approximation error is further exacerbated when using latent diffusion models [61, 38], and can lead to artifacts in image restoration [32].
To address this error source, prior work explores projection-based methods that project either the diffusion state [62] or the measurement gradient [33, 41] onto the data manifold. However, each has limitations: [62] relies on a learned projector at every step, leading to accumulated errors; [33] assumes a noiseless setting; [41] uses SVD, which may not generalize beyond natural images. Other works incorporate higher-order information[63, 64, 65], but most large-scale applications still adopt the isotropic Gaussian approximation for its efficiency, scalability, and simplicity [40].
Our contributions. We propose Equivariance Regularized (EquiReg) diffusion, a symmetry-driven regularization framework for solving inverse problems with diffusion models (Figure 2).
  • Generalized regularization via Wasserstein gradient flow: We reinterpret the reversed conditional diffusion as a Wasserstein-2 gradient flow minimizing a functional over sample trajectories. This framework suggests employing a regularizer that reweights and penalizes trajectories deviating from the data manifold. This addresses errors arising in the likelihood score due to poor approximations of the posterior $p_{t}({\bm{x}}_{0}|{\bm{x}}_{t})$, as well as prior score errors resulting from off-manifold trajectories. Unlike prior works, this approach provides a generalized regularization strategy over the entire reverse diffusion process. To realize such a regularizer, we seek an approach for manifold alignment via global properties of the data distribution. We take equivariance as one such global property that can enforce geometric symmetries.
  • EquiReg with distribution-dependent equivariance errors: We propose a new class of regularizers grounded in distribution-dependent equivariance error, a formalism that quantifies how symmetry violations vary depending on whether samples lie on- or off-manifold. We introduce Manifold-Preferential Equivariant (MPE) functions, which exhibit low equivariance error on the support of the data manifold (in-distribution) and higher error off-manifold (out-of-distribution). Our proposed regularization, EquiReg, is based on this MPE property.
  • Identifying manifold-preferential equivariant functions: We present two strategies for finding MPE functions: a) training induced, where equivariance emerges in encoder-decoder architectures trained with symmetry-preserving augmentations, and b) data inherent, where MPE arises from inherent data symmetries, commonly observed in physical systems.
  • Two practical equivariance regularizers: We develop two flexible plug-and-play losses, Equi and EquiCon (Equi with an implicit manifold constraint), which can be integrated into a variety of pixel- and latent-space diffusion models to improve solving inverse problems. Both regularizations guide the sampling trajectory toward symmetry-preserving regions which often lie close to the data manifold, improving posterior sampling.
  • Improved image restoration performance: We demonstrate that EquiReg significantly improves performance on linear and nonlinear image restoration tasks (Tables 1, 2 and 3). For example, Equi-DPS improves the FID of DPS [32] by $59\%$ on super-resolution (×$\times 4$). On motion deblur, EquiCon-ReSample improves the LPIPS of ReSample [37] by $51\%$. Moreover, EquiReg consistently shows improvement in performance across many noise levels (Figure 5).
  • Improved PDE reconstruction performance: We show that EquiReg enhances performance on PDE solving tasks (Table 4). For instance, Equi-FunDPS reduces the relative $\ell_{2}$ error of FunDPS [56, 57] from $17.16\%$ to $15.91\%$ on Helmholtz inverse problems and from $8.48\%$ to $7.84\%$ on Navier-Stokes inverse problems.
2 Related Works
Diffusion models for generation. Diffusion generative models [27, 28, 66, 67] are state-of-the-art in computer vision for image [68, 69] and video generation [61, 70, 71, 72, 73], with score-based methods [30] being among the most widely used. Recent work has leveraged pre-trained diffusion models for text-guided image generation [74, 75, 76, 77, 78], while others, including variational approaches [79, 80, 81], have employed diffusion models to solve inverse problems by conditioning on measurements [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 51, 52, 58, 59]. Another line of work uses text conditioning to improve inverse problem solutions [82]. The most widely adopted text-to-image diffusion models are latent diffusion models (LDMs) [61], which typically use classifier-free guidance [83]. Large-scale diffusion models for image [84, 85] and video generation [71] are also based on LDMs.
Diffusion models for inverse problems. Solving inverse problems with pre-trained diffusion models requires approximating the intractable likelihood score $\nabla_{{\bm{x}}_{t}}\log{p_{t}({\bm{y}}|{\bm{x}}_{t})}$. Diffusion-based inverse solvers differ in how they approximate the likelihood $p_{t}({\bm{y}}|{\bm{x}}_{t})$ and integrate it with the sampling prior $p_{t}({\bm{x}}_{t})$ [63]. To make the likelihood tractable, a common assumption is an isotropic Gaussian approximation of the posterior distribution $p_{t}({\bm{x}}_{0}|{\bm{x}}_{t})$ [32, 86, 60, 38]. Prior work employs the empirical Bayes estimator (known as Tweedie’s formula) [87, 88, 89] to estimate the posterior expectation. Peng et al. [63] provides a unifying perspective on this approach and proposes a non-isotropic Gaussian approximation, aligning with other works using higher-order information [64, 65].
Equivariant deep learning. Equivariance is a strategy for incorporating symmetries into deep learning [90]. Prior work has applied it to graph networks [91], convolutional networks [92, 93], and Lie groups for modeling dynamical systems [94]. The extension of equivariance to probabilistic symmetries [95] has enabled the development of sample efficient diffusion models [96], with applications in molecular generation [97, 98, 99, 100], autonomous driving [101], robotics [102], crystal structure prediction [103], and audio inverse problems [104].
Prior work has used equivariance as a data prior for solving inverse problems such as super-resolution and deblurring [105], either by training the inverter networks with pre-defined equivariance properties [106], or enforcing it on learned denoisers [107]. Their empirical benefits are supported by theoretical analysis in compressed sensing, showing that equivariance enables solving inverse problems with fewer measurements [108]. In diffusion models, equivariance guidance has been applied to improve temporal consistency in video generation [70], and to break symmetry for generation [109]. These prior studies focus on denoising scores for generation, while we use equivariance to correct likelihood-induced errors in inverse problems and restoration.
3 Preliminaries
@@
Diffusion models. Diffusion models [27, 28, 29, 66, 67] generate data via a reverse noising process. The forward noising process transforms the data sample ${\bm{x}}_{0}\sim p_{\text{data}}$ via a series of additive noise into an approximately Gaussian distribution ($p_{\text{data}}\rightarrow p_{t}\rightarrow\mathcal{N}(0,I)$ as $t\rightarrow\infty$), described by the stochastic differential equation (SDE) [30] $\mathrm{d}{\bm{x}}=-\tfrac{\beta_{t}}{2}{\bm{x}}\mathrm{d}t+\sqrt{\beta_{t}}\mathrm{d}{\bm{w}}$, where $\bm{w}$ is a standard Wiener process, and the drift and diffusion coefficients are parameterized by a monotonically increasing noise scheduler $\beta_{t}\in(0,1)$ in time $t$ [27]. Reversing the forward diffusion process is described by [110]
$\mathrm{d}{\bm{x}}=[-\tfrac{\beta_{t}}{2}{\bm{x}}-\beta_{t}\nabla_{{\bm{x}}_{t}}\log{p_{t}({\bm{x}}_{t})}]\ \mathrm{d}t+\sqrt{\beta_{t}}\mathrm{d}\bm{\bar{w}}$ (2)
@@
with $\mathrm{d}t$ moving backward in time. In discrete steps that would be from $T$ to 00. This reverse SDE can be used to generate (sample) data from the distribution $p_{\text{data}}$, where the unknown gradient $\nabla_{{\bm{x}}_{t}}\log{p_{t}({\bm{x}}_{t})}$ can be approximated by a scoring function $s_{\theta}({\bm{x}}_{t},t)$, parameterized by a neural network, and learned via denoising score matching methods [111, 31]. Solving inverse problems is described as a conditional generation where the data is sampled from the posterior $p({\bm{x}}|{\bm{y}})$, i.e.,
@@
Challenges of training-free diffusion models for inverse problems. Given the intractability of the likelihood, $p_{t}({\bm{y}}|{\bm{x}}_{t})=\int{p({\bm{y}}|{\bm{x}}_{0})p_{t}({\bm{x}}_{0}|{\bm{x}}_{t})\mathrm{d}{\bm{x}}_{0}}$, training-free diffusion models account assumptions on $p_{t}({\bm{x}}_{0}|{\bm{x}}_{t})$. The most popular approximation is the isotropic Gaussian approximation $\mathcal{N}({\bm{x}}_{0|t},r^{2}_{t}{\bm{I}})$ [32, 38, 60, 39, 40]. With an optimal denoising score $s_{\theta}({\bm{x}}_{t},t)$, the posterior expectation ${\bm{x}}_{0|t}\coloneqq\mathbb{E}[{\bm{x}}_{0}|{\bm{x}}_{t}]$ can be computed via Tweedie’s formula. While this offers an MMSE estimation of the posterior expectation, for complex distributions, there is no guarantee that the distribution $p_{t}({\bm{x}}_{0}|{\bm{x}}_{t})$ is concentrated around its mean. From the manifold perspective, this can be interpreted as having a posterior expectation that does not lie on the data manifold $\mathcal{M}$. Error in this estimation can propagate beyond the likelihood score into the whole posterior sampling process. Our proposed EquiReg aims to regularize the process to account for these errors.
Equivariance. An equivariant function respects symmetries under group transformations, i.e.,

### v2

# arXiv:2505.22973v2[cs.LG] 01 Mar 2026

@@ Abstract
Diffusion models represent the state-of-the-art for solving inverse problems such as image restoration tasks. Diffusion-based inverse solvers incorporate a likelihood term to guide prior sampling, generating data consistent with the posterior distribution. However, due to the intractability of the likelihood, most methods rely on isotropic Gaussian approximations, which can push estimates off the data manifold and produce inconsistent, poor reconstructions. We propose Equivariance Regularized (EquiReg) diffusion, a general plug-and-play framework that improves posterior sampling by penalizing trajectories that deviate from the data manifold. EquiReg formalizes manifold-preferential equivariant functions that exhibit low equivariance error for on-manifold samples and high error for off-manifold ones, thereby guiding sampling toward symmetry-preserving regions of the solution space. We highlight that such functions naturally emerge when training non-equivariant models with augmentation or on data with symmetries. EquiReg is particularly effective under reduced sampling and measurement consistency steps, where many methods suffer severe quality degradation. By regularizing trajectories toward the manifold, EquiReg implicitly accelerates convergence and enables high-quality reconstructions. EquiReg consistently improves performance in linear and nonlinear image restoration tasks and solving partial differential equations.
@@ 1 Introduction
Inverse problems aim to recover an unknown signal ${\bm{x}}^{*}\in\mathbb{R}^{d}$ from undersampled noisy measurements:
@@
where $\mathcal{A}$} is a known measurement operator, and $\bm{\nu}$ is an unknown noise (Groetsch, 1993). Inverse problems are widely studied in science and engineering, including imaging and astrophotography.
Inverse problems are ill-posed, i.e., the inversion process can have many solutions; hence, they require prior information about the desired solution (Kabanikhin, 2008). In the Bayesian formulation, the solution maximizes the posterior distribution $p({\bm{x}}|{\bm{y}})\propto p({\bm{y}}|{\bm{x}})p({\bm{x}})$, where $p({\bm{y}}|{\bm{x}})$ is the likelihood of the measurements and $p({\bm{x}})$ is a prior describing the signal structure (Stuart, 2010). Examples of handcrafted priors include sparsity (Donoho, 2006) and low-rankness (Candès et al., 2011).
Figure 1: Equivariance Regularized (EquiReg) diffusion for inverse problems. (left) Manifold preferential equivariance (MPE) functions whose equivariance error is lower for on-manifold and higher for off-manifold data. (right) EquiReg regularizes the posterior sampling trajectory for improved performance. It penalizes off-manifold trajectories via MPE-based regularization.
Figure 2: EquiReg’s broad applicability. a-d) image restoration inverse problems and e) text-guided image generation, resulting in artifact reduction and more realistic generation. Here, EquiReg refers to our regularization being applied to the diffusion sampling method on the same row.
@@
This paper focuses on methods that leverage unconditionally pre-trained score‐based generative diffusion models as learned priors (Ho et al., 2020; Song and Ermon, 2019) with applications in image restoration (Chung et al., 2023), medical imaging (Chung et al., 2022a), and solving partial differential equations (PDEs) (Huang et al., 2024; Yao et al., 2025). These methods define a sequential noising process ${\bm{x}}_{0}\sim p_{\mathrm{data}}\to{\bm{x}}_{t}\to{\bm{x}}_{T}\sim p_{T}({\bm{x}})\approx\mathcal{N}(\bm{0},{\bm{I}})$ and a reverse denoising process parameterized by a neural network score $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{x}}_{t})$ (Vincent, 2011). During sampling, these approaches incorporate gradient signals carrying likelihood information to solve inverse problems.
@@
Solving inverse problems with diffusion (Zhang et al., 2025a; Alkhouri et al., 2025) requires computing the conditional score $\nabla_{{\bm{x}}_{t}}\log{p_{t}({\bm{x}}_{t}|{\bm{y}})}$, decomposed into $\nabla_{{\bm{x}}_{t}}\log{p_{t}({\bm{x}}_{t})}+\nabla_{{\bm{x}}_{t}}\log{p_{t}({\bm{y}}|{\bm{x}}_{t})}$. This introduces challenges, as the likelihood score $\nabla_{{\bm{x}}_{t}}\log{p_{t}({\bm{y}}|{\bm{x}}_{t})}=\nabla_{{\bm{x}}_{t}}\log{\int{p({\bm{y}}|{\bm{x}}_{0})p_{t}({\bm{x}}_{0}|{\bm{x}}_{t})\mathrm{d}{\bm{x}}_{0}}}$ is only computationally tractable when $t=0$. To handle the likelihood for $t>0$, many methods approximate the posterior $p_{t}({\bm{x}}_{0}|{\bm{x}}_{t})$ with the isotropic Gaussian distribution (Zhang et al., 2025a), where the distribution expectation is computed using the optimal denoising score (Robbins, 1956). The Gaussian approximation can be inaccurate for complex distributions (Figure˜3), leading to errors in likelihood computation, especially with point estimations (Chung et al., 2023). Since the posterior expectation is a conditional expectation, a linear combination of all possible ${\bm{x}}_{0}$, it may lie off the data manifold even when individual samples remain on it. These issues are further amplified in latent diffusion models (LDMs), introducing artifacts (Rout et al., 2023).
Prior work has attempted to address this challenge via projection-based (He et al., 2024; Zirvi et al., 2025) or decoupled optimization strategies (Zhang et al., 2025a), aimed at reducing the propagation of measurement errors during sampling. However, they rely on the isotropic Gaussian assumption, which can lead to failures on difficult tasks or at reduced sampling steps. While higher-order statistics reduces errors (Boys et al., 2024), most approaches employ such approximation for its efficiency, scalability, and simplicity (Alkhouri et al., 2025), often coupled with large-scale LDMs (Peebles and Xie, 2023). This raises a key question: how can we ensure the reliability and practicality of conditional diffusion models under this approximation?
Equivariance offers a natural mechanism to keep sampling trajectories close to the data manifold. We address this challenge with a regularization scheme that leverages equivariance to improve posterior sampling by guiding trajectories toward symmetry-preserving solution spaces. Prior work has enforced equivariance directly within generation or denoising processes (Chen et al., 2023a; Terris et al., 2024), with extensions to probabilistic symmetries (Bloem-Reddy et al., 2020) enabling sample efficiency (Wang et al., 2024).
Our approach differs as follows: rather than strictly enforcing equivariance within denoising architectures, which can hinder tasks requiring symmetry breaking (Lawrence et al., 2025), we employ equivariance as a plug-and-play regularizer to guide diffusion trajectories toward the data manifold.
Our contributions. We propose Equivariance Regularized (EquiReg) diffusion, an equivariance-based regularization framework for solving inverse problems with diffusion models (Figure˜1). EquiReg leverages equivariance to regularize likelihood-induced errors during posterior sampling, guiding diffusion trajectories toward more consistent, on-manifold solutions. Crucially, it employs Manifold-Preferential Equivariant (MPE) functions, which discriminate on-manifold from off-manifold data by exhibiting low equivariance error in-distribution and higher error out-of-distribution.
Figure 3: Off-manifold posterior expectation. This impacts the likelihood score $p_{t}({\bm{y}}|{\bm{x}}_{t})=\int{p({\bm{y}}|{\bm{x}}_{0})p_{t}({\bm{x}}_{0}|{\bm{x}}_{t})\mathrm{d}{\bm{x}}_{0}}$ computation achieved via isotropic Gaussian modelling of $p_{t}({\bm{x}}_{0}|{\bm{x}}_{t})$.
We formalize that an effective regularizer should capture such a global property, and MPE functions provide a principled way to direct sampling toward plausible solutions. This design makes EquiReg architecture-agnostic: the regularizer operates independently of the diffusion model itself. With a suitable MPE function, EquiReg improves performance across models, including those with equivariant scores, where likelihood guidance may otherwise push trajectories off the manifold.
We observe that many practical functions behave as MPEs: their equivariance error is small on the training or data manifold but grows off-manifold. This behavior arises in learned models trained with data augmentation, as well as in data with inherent symmetries such as those from physical systems. Rather than treating the degradation off-manifold as a limitation, we exploit it as a signal: equivariance error serves as a natural discriminator for identifying undesirable states during diffusion sampling. Building on this idea, we construct pre-trained MPEs as the foundation of our EquiReg loss. The choice of this function is independent of the denoiser in diffusion models and can be derived separately. For instance, if the diffusion architecture is itself equivariant to a set of group actions, it cannot be regularized via EquiReg using these same equivariances, as the MPE cannot discriminate between on- and off-manifold samples (the equivariance loss would remain low). In this case, a separate MPE with different equivariance properties can be trained. We explore several constructed and pre-trained MPE functions and show that EquiReg is widely robust to the choice of MPE.
Beyond architectural flexibility, we systematically analyze the diversity properties of EquiReg-guided posterior sampling. We demonstrate that EquiReg offers favourable fidelity–diversity trade-offs, and as inverse problems become more ill-posed, diversity increases approximately linearly, indicating that EquiReg naturally expands posterior exploration with growing uncertainty rather than collapsing to a single mode.
We further demonstrate that MPE functions are not rare or specialized constructs, but emerge broadly across practical neural networks. We empirically analyze multiple architectures, including the latent diffusion encoder (Rombach et al., 2022), CNN autoencoders trained with symmetry-based augmentations, pre-trained ResNet-50 (He et al., 2016), and CLIP encoders (Radford et al., 2021), and show that their equivariance error increases systematically as Gaussian noise pushes samples off the data manifold. We further show that using these MPE functions for EquiReg guidance improves solutions to inverse problems. Overall, this study shows that MPEs are widespread, and EquiReg leverages this intrinsic property as a principled regularization signal.
We validate the efficacy of EquiReg through extensive experiments across diverse diffusion models, inverse problems, and datasets. We demonstrate that EquiReg improves perceptual image quality and remains effective in cases where baselines fail. We show that EquiReg improves the performance of SITCOM (Alkhouri et al., 2025) and DPS (Chung et al., 2023) when the number of measurement consistency and sampling steps are reduced, thus moving toward more efficient diffusion-based solvers. Our method is particularly useful when applied to LDMs. EquiReg reduces failure cases, and consistently improves PSLD (Rout et al., 2023), ReSample (Song et al., 2023a), and DPS (Chung et al., 2023) on linear and nonlinear image restoration tasks. For example, EquiReg significantly improves the LPIPS (Song et al., 2023a) of ReSample by $51\%$ for motion deblur and the FID of DPS (Chung et al., 2023) by $59\%$ on super-resolution. We also include diversity analyses, demonstrating that EquiReg maintains diversity without collapse of single mode reconstruction.
We extend EquiReg’s applicability to function-space diffusion models and demonstrate its added benefit for solving PDEs. EquiReg achieves a $7.3\%$ relative reduction in the $\ell_{2}$ error of FunDPS (Mammadov et al., 2024a; Yao et al., 2025) on the Helmholtz equation and a $7.5\%$ relative reduction on the Navier-Stokes equation. Lastly, we include preliminary experiments on EquiReg improving the realism and plausibility of text-guided image generation, emphasizing that the benefits of EquiReg extend beyond image restorations. Overall, the flexibility of EquiReg as a plug-and-play regularization framework suggests that its utility will extend well beyond the specific methods studied in this paper.
2 Preliminaries and Related Works
@@
Diffusion models. Diffusion generative models (Ho et al., 2020; Song and Ermon, 2019; Sohl-Dickstein et al., 2015; Kadkhodaie and Simoncelli, 2021) are state-of-the-art in computer vision for image (Esser et al., 2024) and video generation (Brooks et al., 2024; Zhang et al., 2025b), with score-based methods (Song et al., 2021) being among the most widely used. Diffusion models generate data via a reverse noising process. The forward noising process transforms the data sample ${\bm{x}}_{0}\sim p_{\text{data}}$ via a series of additive noise into an approximately Gaussian distribution ($p_{\text{data}}\rightarrow p_{t}\rightarrow\mathcal{N}(0,I)$ as $t\rightarrow\infty$), described by the stochastic differential equation (SDE) $\mathrm{d}{\bm{x}}=-\tfrac{\beta_{t}}{2}{\bm{x}}\mathrm{d}t+\sqrt{\beta_{t}}\mathrm{d}{\bm{w}}$, where $\bm{w}$ is a standard Wiener process, and the drift and diffusion coefficients are parameterized by a monotonically increasing noise scheduler $\beta_{t}\in(0,1)$ in time $t$ (Ho et al., 2020). Reversing the forward diffusion process is described by (Anderson, 1982)
$\vskip-5.69054pt\mathrm{d}{\bm{x}}=[-\tfrac{\beta_{t}}{2}{\bm{x}}-\beta_{t}\nabla_{{\bm{x}}_{t}}\log{p_{t}({\bm{x}}_{t})}]\ \mathrm{d}t+\sqrt{\beta_{t}}\mathrm{d}\bm{\bar{w}}$ (2)
@@
with $\mathrm{d}t$ moving backward in time or in discrete steps from $T$ to 0. This reverse SDE is used to sample data from the distribution $p_{\text{data}}$, where the unknown gradient $\nabla_{{\bm{x}}_{t}}\log{p_{t}({\bm{x}}_{t})}$ is approximated by a scoring function $s_{\theta}({\bm{x}}_{t},t)$, parameterized by a neural network and learned via denoising score matching methods (Hyvärinen and Dayan, 2005; Vincent, 2011). Solving inverse problems is described as a conditional generation where the data is sampled from the posterior $p({\bm{x}}|{\bm{y}})$:
@@
Diffusion models for inverse problems. Solving inverse problems with pre-trained diffusion models requires approximating the intractable likelihood score $\nabla_{{\bm{x}}_{t}}\log{p_{t}({\bm{y}}|{\bm{x}}_{t})}$. Training-free solvers differ in how they approximate $p_{t}({\bm{y}}|{\bm{x}}_{t})$ and combine it with the prior $p_{t}({\bm{x}}_{t})$ (Peng et al., 2024). Since $p_{t}({\bm{y}}|{\bm{x}}_{t})=\int{p({\bm{y}}|{\bm{x}}_{0})p_{t}({\bm{x}}_{0}|{\bm{x}}_{t})\mathrm{d}{\bm{x}}_{0}}$, the common choice is to approximate $p_{t}({\bm{x}}_{0}|{\bm{x}}_{t})$ by an isotropic Gaussian $\mathcal{N}({\bm{x}}_{0|t},r_{t}^{2}{\bm{I}})$ (Chung et al., 2023; Song et al., 2023b; Zhu et al., 2023; Zhang et al., 2025a). With an optimal denoising score $s_{\theta}({\bm{x}}_{t},t)$, the posterior mean ${\bm{x}}_{0|t}\coloneqq\mathbb{E}[{\bm{x}}_{0}|{\bm{x}}_{t}]$ follows from Tweedie’s formula (Robbins, 1956; Miyasawa and others, 1961; Efron, 2011). While an MMSE estimate, $p_{t}({\bm{x}}_{0}|{\bm{x}}_{t})$ may not be concentrated around its mean for complex or multimodal distributions, leading to off-manifold solutions (see Figure˜3).
Equivariance. Equivariance describes how functions transform predictably under group actions and provides a principled way to incorporate symmetry into deep learning (Bronstein et al., 2021). It has been applied to graphs (Satorras et al., 2021), convolutional networks (Cohen and Welling, 2016; Romero and Lohit, 2022), Lie groups for dynamical systems (Finzi et al., 2020), and diffusion models (Wang et al., 2024), with applications spanning molecular generation (Hoogeboom et al., 2022; Cornet et al., 2024), autonomous driving (Chen et al., 2023b), robotics (Brehmer et al., 2023), crystal structure prediction (Jiao et al., 2023), and audio inverse problems (Moliner et al., 2023). Equivariance guidance has also improved temporal consistency in video generation (Daras et al., 2024), and its role as a prior in inverse problems is theoretically supported in compressed sensing (Tachella et al., 2023).